# Aharonov-Bohm oscillations and perfectly transmitted mode in amorphous topological insulator nanowires

### v1

@@ VI Discussion
Using local topological markers, we have seen that the conductance resonances act as transport signatures of a phase transition away from a topological phase. In particular, amorphicity drives the nanowires from a topological phase characterized by a quantized chiral marker $\nu=-1$ to an insulating phase where the topology is not well defined. Although our calculations in the fully amorphous case focus on the chiral limit, we argued that our results, both in transport and topological characterization, are representative for the regime where the perfectly transmitted mode is the most relevant contribution to conductance, even in the absence of chiral symmetry.
Our results contribute to understanding how transport and local topological markers can be used as a probe for amorphous topological states, and explain how certain amorphous structures and symmetries, including statistical symmetries Fulga et al. (2014); Corbae et al. (2023b); Spring et al. (2021, 2023), may be important in order to create amorphous topological nanowires for further applications. Moreover, the nanowires we have used are also accessible in experiments, since amorphous nanowires with a smooth and approximately constant cross section may be achieved by FIB etching a nanowire Hunter et al. (2024); Estry et al. (2024). The methods we used could also be potentially useful to explain the transport properties of other amorphous materials, such as for example the anomalous resistivity scaling with thickness seen in thin films of amorphous CoSi Rocchino et al. (2024).

### v2

@@ VI Discussion
Using local topological markers, we have seen that the conductance resonances act as transport signatures of a phase transition away from a topological phase. In particular, amorphicity drives the nanowires from a topological phase characterized by a quantized chiral marker $\nu=-1$ to an insulating phase where the topology may not be well defined. Although our calculations in the fully amorphous case focus on the chiral limit, we argued that our results, both in transport and topological characterization, are representative for the regime where the perfectly transmitted mode is the most relevant contribution to conductance, even in the absence of chiral symmetry.
With the parameters chosen in our calculations, the clean bulk band gap is of the order of $\Delta E\sim 2t$, with a Fermi velocity of $v_{F}\sim 1.7$, giving a localization length for the surface states in a clean nanowires of $\xi=v_{F}/\Delta E\sim 1$ in units of the lattice spacing. In our calculations $N$ is sufficiently large compared to $\xi$ for finite size effects to be suppressed in the clean and low amorphicity regime. For the topological transition to occur, the mobility gap should close increasing the localization length to values larger than the cross section. Hence, at large amorphicities the marker may not accurately capture the topological phase depending on the chosen parameters, explaining the lack of quantization of the marker at high amorphicity.
The nanowires we have used are also accessible in experiments, since amorphous nanowires with a smooth and approximately constant cross section may be achieved by FIB etching a nanowire [106; 107]. Experimentally, the magnetic field needed to see the phenomena we describe is the same for crystalline and amorphous nanowires, and depends on the cross section size as the flux needs to be of the order of one flux quanta. The dimensions of experimental nanowires vary depending on the particular work, but their perimeter ranges usually from 100450100-450nm such that the magnetic field required to get $\phi=h/e$ ranges from 0.3100.3-10T [19]. Although in our calculations we cannot reach experimentally relevant sizes, the behavior we see is relevant for experiments as long as the level broadening is smaller than the level spacing, as is our case.
Our results contribute to understanding how transport and local topological markers can be used as a probe for amorphous topological states, and explain how certain amorphous structures and symmetries, including statistical symmetries [96; 62; 94; 95], may be important in order to create amorphous topological nanowires for further applications. The methods we used could also be potentially useful to explain the transport properties of other amorphous materials, such as for example the anomalous resistivity scaling with thickness seen in thin films of amorphous CoSi [29].